OpenStudy (anonymous):

how do i solve the following: 3/4X-2/3X=2 -8X+9+6X=-3X+14 1/3X+1/6X=2/5X-7/10 X+14-6X=2X-7 -2(X-1)-3(2X+1)=-9

OpenStudy (anonymous):

isolate the variable by adding, subtracting, dividing and multiplying, but I recommend multiplying out your fractions first. ALso you will probably have to use the distributive property on the last question.

OpenStudy (anonymous):

3/4x-2/3x=2 3x-2x=24 combine like terms x=24 18-16=2

OpenStudy (anonymous):

I TRIED BUT AM NOT GETTING CORRECT ANSWER

OpenStudy (anonymous):

-8X+9+6X=-3X+14 -8x+6x+3x=14-9 x=5

OpenStudy (anonymous):

X+14-6X=2X-7 x-6x-2x=-14-7 -7x=-21 x=3

OpenStudy (anonymous):

THANKS 2 MORE QUESTIONS? 1/3X+1/6=2/5X-7/10 -2(X-1)-3(2X+1)=-9